L020 composite depreciation ur decimals (based on cost) and the composite life. California Utilities owns a power plant that Exercise 12-59 Recording Entries Using Composite consists of the following assets, all acquired on January 1, 2020. Cost Estimated Residual Value Estimated Life (Years) Depreciation Method L03 300,000 $10,000 0 5,000 Machinery 95,000 50,000 30 10 Required a. Compute the total straight-line depreciation for 2020 on all items combined. b. Compute the composite depreciation rate rounded to two decimals (based on cost) and the composite life of the plant. c. Provide the entry to record 2020 composite depreciation d. Assume that all of the Other equipment was sold in 2023 for $6,000 cash. Prepare the entry for the sale of that equipment.
Answer #1
LO3
a: Total Straight line depreciation:
COST SALVAGE DEPRECIABLE VALUE EST. LIFE DEP. PER YEAR
BULD 300000 10000 290000 30 9666.67
MACHI 95000 0 95000 10 9500
OTHER EQ. 50000 5000 45000 6 7500
TOTAL DEPRECIATION 26666.67
b)
COST SALVAGE DEPRECIABLE VALUE EST. LIFE DEP. PER YEAR
BULD 300000 10000 290000 30 9666.67
MACHI 95000 0 95000 10 9500
OTHER EQ. 50000 5000 45000 6 7500
TOTAL 445000.00 15000.00 430000.00 26666.67
COMPOSITE RATE = Total depreciation /Total cost = 26666.67/445000 = 0.0599
COMPOSITE LIFE = Total Depreciable cost / total depreciation per year = 430000/26666.67=16.125 years
c) journal entry:
Accounts Titles Debit Credit
Debit Depreciation Exp 26666.67
credit Acc. Dep. 26666.67
(being composite depreciation recorded)
d)
Debit Acc. Dep 30000 7500*4
cash 6000
Loss on Sale 14000
Other Equip 50000
(being sale of other equip in 2023)